On September 30, 2019, Morgan, Inc., acquired all of the outstanding common stock of Pathways, Inc., for $100 million. In addition to tangible assets, Morgan recorded the following assets as a result of the acquisition: Patent $6 million Developed technology 3 million In-process research & development 2 million Goodwill 7 million Morgan's policy is to amortize intangible assets using the straight-line method, no residual value and a six-year useful life. Required: Record the journal entry to record amortization of the above intangible assets.
Answer #1

Patent:$6m/6=$1m

Developed technology:$3 m /6=$0.5m

In process research & development: $2m/6=0.33m

Good will: $7m/6=$1.167

Journal entries:

Amortization a/c Dr $2.997m

To Patent a/c $1m

To developed technology a/c. $0.5m

To research & development a/c. $0.33m

To gooe will a/c. $1.167

(Being amortization of intangible assets over a period of six years )

Note:( Every year for a period of six years same entry)